I have to use two different types of milk (eg. Milk A & B)which have two different protein to fat ratios - in percents per 100 litre of milk.
In milk A, the protein to fat ratio is 3.22/ 3.85. = .836
In milk B, the protein to fat ratio is 5.18/ 2.67 = 1.94
What quantities of milk A and B should be used to get 7000 L of mix at a ratio of (protein/fat) = 1.01
If possible let me know the equation used for this
thanks
However I tried with the eqn
A+B = 7000, so A = 7000-B
A(.836) + B(1.94) = 7000(1.01)
I got the answer also but When I check their Quantities multiply with their Individual Protein and Fat ; then their ratios are not matching
e.g 5899 Qty of milk Fat (A) + 1101 Qty of milk Fat (B) when added up to Total fat and same way for Protein ....and now when we take protein/Fat then it shows diff ratio.....Can U explain this

Answer by jorel1380(3719)

7000x1.01=7070
m+n=7000
.836m+1.94n=7070
836m+1940n=7070000
836m+836n=5852000
1104n=1218000
n=1103.261
m=5896.74
You need 5896.74 liters of .836 milk and 1103.26 liters of 1.94 milk to get your mix..